# Complementarity in Human-AI Collaboration: Concept, Sources, and Evidence

### v1

@@ 4.2.2 Study Design
We conduct an online experiment, employing a between-subjects design with two conditions. Participants are recruited from prolific.com and randomly assigned to one of the conditions; repeated participation is not allowed. We employ a similar experimental set up as in the first experiment (see Appendix C for additional details on the study design):
Step 1: The participants are transferred to the experimental website after accepting the task, after which they have to submit a consent form and answer an initial control question. Thereafter they have to pass an attention check.
Step 2: They are given a task tutorial which presents an exemplary image along with the 16 classes arranged in a four-by-four matrix, including the 16 class icons with their name displayed underneath.
Step 3: The participants are given an introduction to the AI and its decision regarding an exemplary image followed by a control question to verify their understanding. Figure 10 displays the interface that the participants see throughout the experiment when collaborating with the AI. In this experiment, the participants are only shown the AI’s decision, since confidence information generated by two separately trained AI models would introduce a confounder.
@@
Step 4: Next, participants start a practice round comprising three images that need to be classified without AI support in order to familiarize themselves with the classification task.
Step 5: After being informed about the start of the main task, each participant classifies 15 images in a randomized order. The participants provide their own classification for each image, then receive the AI recommendation and are asked to verify and adjust, if required, the AI classification in the best way possible. During the task, they are not informed about the true class of any image. After classifying all the images, participants are asked to answer a questionnaire regarding qualitative feedback (step 6) and demographic information (step 7).
The overall task lasts approximately 20 minutes. Before recruiting participants, we compute the required sample size in a power analysis using G*Power (Faul et al.,, 2007). We test for a medium to large effect ($d=$ 0.65) and refer to an alpha value of 0.05, while taking multiple testing into account in order to achieve a power of 0.8. This results in a total sample size of 128. In order to buffer for participants potentially failing attention checks, we recruit a total of 170 participants. They receive a base payment of £8 and are additionally incentivized following the approach pursued in the first behavioral experiment (Kvaløy et al.,, 2015). We exclude participants who did not pass the integrated attention checks, resulting in 144 participants—76 in the base condition (baseline AI) and 68 in the intervention (complementary AI). We provide further details about the participants in Appendix C.
@@ 4.2.3 Evaluation Measures
For each participant, the loss of the human ($l_{H}$), the AI ($l_{AI}$), and the team decision ($l_{I}$) is measured as the classification error and averaged over all the task instances, providing the human ($L_{H}$), the AI ($L_{AI}$), and the team performance ($L_{I}$). We also calculate the complementarity potential and effect, including their components (see Section 3). Finally, for each measure, we calculate the average over all the participants.
@@ 4.2.4 Results
Figure 11: Performance results as the classification error across the conditions, including 95% confidence intervals. The red horizontal line denotes the AI performance.
@@
We observe a significant increase in the inherent complementarity potential (${CP}^{inh}$) in the condition with the complementary AI (Baseline AI: 0.0640, Complementary AI: 0.2480; $d=$ 3.81, $p<$ 0.001, two-tailed Mann-Whitney U test). This reflects the AI model’s more asymmetric capabilities compared to that of the humans in this condition. Compared to the baseline AI condition, the AI model makes more erroneous decisions for instances that humans can process correctly, whereas humans err for instances that AI can classify correctly. Conversely, we observe a significant decrease in the collaborative complementarity potential (${CP}^{coll}$) in the condition with the complementary AI (Baseline AI: 0.2026, Complementary AI: 0.0186, $d=$ 3.81, $p<$ 0.001, two-tailed Mann-Whitney U test) as the $CP$ remains constant due to the AI being individually more accurate than the humans. Whereas the inherent complementarity potential constitutes 24% of the overall complementarity potential in the baseline condition, this share rises to 93% in the complementary AI condition. This means that, for the majority of the task instances, one team member is theoretically capable of making a correct decision, provided that the human responsible for the team decision relies on his/her own assessment or that of the AI in respect of the correct instances.
@@
In the baseline condition, 58% of the inherent complementarity potential ($\frac{{CE}^{inh}}{{CP}^{inh}}$) could be realized by the humans integrating their own decision and that of the AI into a final team decision, resulting in a ${CE}^{inh}$ of 0.0368. In the condition with the complementary AI, it is possible to realize 89% of the inherent complementarity potential, resulting in a ${CE}^{inh}$ of 0.2196. This not only means a significant performance improvement ($d=$ 3.43, $p<$ 0.001, two-tailed Mann-Whitney U test), but also that a significantly larger fraction of $\frac{{CE}^{inh}}{{CP}^{inh}}$ could be realized ($d=$ 1.02, $p<$ 0.001, two-tailed Mann-Whitney U test). It indicates that humans tended to rely on the AI decisions when they were correct, but on their decision when it was incorrect. Moreover, in both conditions the collaborative complementarity effect (${CE}^{coll}$) is negative. Whereas the value is only slightly negative in the baseline condition (Baseline AI: -0.0175), it decreases to -0.0990 in the condition with the complementary AI. The difference between the two conditions is significant ($d=$ 1.32, $p<$ 0.001, two-tailed Mann-Whitney U test).
Figure 12: Summary of the results of the image classification experiment.
The observed changes in both of the components reveal interesting insights. The increase in the inherent complementarity effect reveals that humans continue to regard the AI as a team member on whom they can rely, especially with regard to the instances that they find difficult, even though they have observed the AI makes several incorrect decisions in respect of the instances that they find easier. Conversely, the decrease in the collaborative complementarity effect reveals that there are also instances where humans, incorrectly, do not rely on the AI. This is an interesting finding, since it could have been conceivable for humans to avoid AI suggestions to a greater extent after witnessing incorrect AI decisions (Dietvorst et al.,, 2015). Consequently, in the team setting in which the human integrates his or her own decision and that of the AI into a final team decision, there might be a trade-off between turning the AI’s more asymmetric capabilities into performance synergies and not relying on its suggestions due to witnessing erroneous decisions in “easier” instances. See Appendix C for additional results, including a performance and a switch fraction analysis of each image. In summary, ${CE}^{inh}$ and ${CE}^{coll}$ result in the total complementarity effect ($CE$)—equivalent to the performance difference between the best individual team member and the team performance (Baseline AI: 0.0193, Complementary AI: 0.1206). Figure 12 summarizes the analysis.
@@ 5 Discussion
Our research emphasizes the importance of considering $CTP$ in human-AI decision-making. In the past, when early AI-based systems still had limited capabilities, $CTP$ was rarely a focus, as these systems could only provide humans with (partial) decision support within a small range of low-stakes decision-making tasks (e.g., calculating decision-making inputs) (Turban et al.,, 2011). Today, however, AI can perform a growing number of tasks independently with human-level performance, and is increasingly utilized in high-stakes decision-making domains such as medicine (McKinney et al.,, 2020), law (Hillman,, 2019), and finance (Day et al.,, 2018). It can also provide general-purpose support, driven by recent advances in large language models that enable applications such as ChatGPT (Bubeck et al.,, 2023). While this opens up the potential for task automation, it also enables a new form of human-AI decision-making with AI models and humans being equitable team members.
@@ 5.1 Contributions
Current research lacks a concise concept of human-AI complementarity, as the majority of studies focusing on human-AI decision-making do not to achieve $CTP$ (Bansal et al.,, 2021; Hemmer et al.,, 2021). Moreover, merely observing whether $CTP$ has been achieved or not, does not allow in-depth conclusions to be drawn about the synergetic potential of humans and AI through collaboration. In this work, we contribute to a comprehensive understanding of human-AI teams’ inner workings regarding decision-making, and provide guidance on how to achieve $CTP$ more consistently.
@@
More specifically, we develop a conceptualization of human-AI complementarity that introduces the notion of complementarity potential in a formalization and delineates relevant sources. This conceptualization allows for analyzing human-AI teams’ potential for decision-making synergies and for making this potential measurable. By differentiating between the inherent and collaborative components of complementarity potential and its realization, it is also possible to gain important insights into the collaboration’s functioning in terms of the optimal joint team decision. These insights, along with the sources of complementarity potential, could inform human-AI teams’ future designs and collaboration mechanisms, depending on the use case and application domain. Furthermore, we not only demonstrate the conceptualization’s utility in two behavioral experiments with humans in the role of integrating the final team decision, but we also empirically show that both information and capability asymmetry can lead to $CTP$. In this context, our conceptualization allows us to reveal interesting empirical insights. The first experiment highlights that providing humans with unique contextual information not only affects the inherent complementarity potential, but can also increase the realized amount, i.e., the inherent complementarity effect, disproportionally. This means that the effectiveness of the integration that humans conducted improved, which is an interesting finding. Intuitively, the perception of having more information than the other team member could alternatively result in reduced utilization of the AI suggestions in the team prediction—resulting in a sub-optimal team performance (Jussupow et al.,, 2020; Mahmud et al.,, 2022; Sieck and Arkes,, 2005). The second experiment reveals that if the human-AI team’s capability asymmetry is greater, this can contribute to the realization of a larger share of the existing complementarity potential. This is also an insightful observation, since humans observing erroneous AI decisions for instances that they could solve relatively easily themselves, might lead them to refute AI support even in situations when it is actually helpful (Dietvorst et al.,, 2015). This finding also contributes to our understanding of team diversity’s compositional impact on performance (Horwitz,, 2005). Whereas expertise diversity in human teams could be a performance driver, e.g., because it fosters a broader range of cognitive skills (Cohen and Bailey,, 1997), it could also turn into a performance inhibitor, e.g., due to the potential difficulties of achieving a mutual understanding (Dougherty,, 1992). As in human teams, there might also be a trade-off, since humans could start avoiding AI suggestions if they observe them erring too often (Dietvorst et al.,, 2015). Our conceptualization enables an analysis of this tradeoff—finding that, in the specific experiment, the capability asymmetry between humans and AI affects team performance positively.
@@ 5.2 Theoretical Implications
From a theoretical perspective, the proposed conceptualization provides a foundation for future research in human-AI decision-making. We offer the research community concrete measures that allow the investigation of human-AI decision-making’s inner workings at a deeper level than aggregate performance comparisons in behavioral experiments allow. In this context, these measures could also help researchers formulate and test hypotheses about the effect of different socio-technical factors on the realization of decision-making synergies in human-AI teams (Jain et al.,, 2021). Our research’s most important implication is the need to design for $CTP$, which is influenced by the complementarity potential and the collaboration mechanism to derive a team decision. Both could and should be purposefully designed. The inherent complementarity potential can be influenced by increasing unique knowledge. From an AI perspective, this could be achieved by deliberately creating “complementary” AI models designed for team work that perform well in areas of the feature space where humans experience difficulties (Hemmer et al.,, 2022; Mozannar and Sontag,, 2020; Wilder et al.,, 2020). From a human perspective, humans could be trained to focus on their unique capabilities and to build awareness to use unique contextual information. Lastly, realizing complementarity potential also implies that the collaboration mechanism should be consciously designed to optimize the realization of inherent and collaborative complementarity potential.
@@ 5.3 Managerial Implications
Our work also has important implications for managerial decision-makers. In application areas suitable for human-AI decision-making tasks, responsible managers should focus on deploying AI systems that enable the realization of $CTP$. If they don’t, their competitors could gain advantages. They could start by collecting data to train AI models that are compatible with team work and to invest in trainings to upskill their employees in order to enhance their human-AI teams’ inherent complementarity potential. The conceptualization supports the identification of use cases suitable for human-AI decision-making and their evaluation of $CTP$’s realization. Current AI endeavors often result in the blind adoption of human-AI collaboration due to decision-makers’ concerns about full automation (Jussupow et al.,, 2021). However, it is worth designing human-AI collaboration purposefully, as this could contribute to a reduction in erroneous decisions with potentially high costs. Rather than fearing automation, decision-makers should explore the benefits of working with AI. This is an important perspective, also in the broader societal debate on the role of AI in automation and in the future of work. Our research provides managers with valuable guidance by helping them determine when and how to collaborate with AI to optimize decision-making and achieve $CTP$.
@@ 5.4 Limitations
Our current research has several limitations that future work needs to address. First, the prevalent use of laboratory-based experiments and vignette methodologies in extant literature on human-AI collaboration, including our own, could limit our work’s generalizability and its practical implications. This underscores the need for future studies to use a nuanced approach to prevent contributing to the fragmentation of research in this domain. Furthermore, we focus on developing measurement tools, delineating sources of complementarity potential, and on validating the proposed concepts in behavioral experiments. While our experiments’ controlled settings allow us to derive the insights presented in this work, they do not yet address the wide range of complex human factors, such as motivation (Schunk,, 1995), engagement (Chandra et al.,, 2022), or behavior patterns within teams (Schecter et al.,, 2022), which also contribute to human-AI collaboration’s effectiveness (Chandra et al.,, 2022). In this context, many factors in socio-technical systems are highly interconnected and influence the underlying systems’ acceptance and use (Jain et al.,, 2021). While a broader exploration of these factors is beyond the scope of our work, advancing the field of human-AI collaboration further also requires their incorporation. We hope that our research will serve as a foundation of studies aiming at expanding knowledge of the mechanisms governing human-AI collaboration.
Another limitation is the way in which we measure the counterfactual human decision if the user has not received advice on using AI. In this work, we used a sequential decision-making setup to first measure the human decision and, thereafter, the team decision. However, the sequential nature of the decision-making process could also influence human behavior. Consequently, the timing when the AI’s recommendation is revealed is another critical aspect (Jussupow et al.,, 2021). Giving the participants the AI’s recommendation upfront, could lead to cognitive capacity being less invested in the task, because the AI has already provided a possible answer (Green and Chen,, 2019).

### v2

@@ 4.2.2 Study Design
We conducted an online experiment, employing a between-subject design with two conditions. Participants were recruited from prolific.com and randomly assigned to one of the conditions; repeated participation was not allowed. We employed a similar experimental set up as in the first experiment (see Figure C.1 in Appendix C.2 for additional details):
Step 1: The participants were transferred to the experimental website after accepting the task, after which they had to submit a consent form and answer an initial control question. Thereafter they had to pass an attention check.
Step 2: They were given a task tutorial which presents an exemplary image along with the 16 classes arranged in a four-by-four matrix, including the 16 class icons with their name displayed underneath (see Figure C.2 in Appendix C.2).
Step 3: The participants were given an introduction to the AI and its decision regarding an exemplary image followed by a control question to verify their understanding (see Figure C.3 in Appendix C.2). Figure 10 displays the interface that the participants see throughout the experiment when collaborating with the AI. In this experiment, the participants were only shown the AI’s decision, since confidence information generated by two separately trained AI models would have introduced a confounder.
@@
Step 4: Next, participants started a practice round comprising three images that had to be classified without AI support in order to familiarize themselves with the classification task (see Figure C.4 in Appendix C.2 for an exemplary training task instance).
Step 5: After being informed about the start of the main task, each participant classified 15 images in a randomized order (see Figure C.5 in Appendix C.2 for an exemplary task instance). The participants provided their own classification for each image, then received the AI recommendation and were asked to verify and adjust, if required, the AI classification in the best way possible. During the task, they were not informed about the true class of any image. After classifying all the images, participants were asked to answer a questionnaire regarding qualitative feedback (Step 6) and demographic information (Step 7).
The overall task lasted approximately 20 minutes. Before recruiting participants, we computed the required sample size in a power analysis using G*Power (Faul et al., 2007). We tested for a medium to large effect ($d=$ 0.65) and considered an alpha value of 0.05, while taking multiple testing into account in order to achieve a power of 0.8. This resulted in a total sample size of 128. In order to buffer for participants potentially failing attention checks, we recruited a total of 170 participants. They received a base payment of £8 and were additionally incentivized following the approach pursued in the first behavioral experiment (Kvaløy et al., 2015). We excluded participants who did not pass the integrated attention checks, resulting in 144 participants—76 in the base condition (baseline AI) and 68 in the intervention (complementary AI). We provide further details in Table C.1 in Appendix C.3.
@@ 4.2.3 Evaluation Measures
For each participant, the loss of the human ($l_{H}$), the AI ($l_{AI}$), and the team decision ($l_{I}$) was measured as the classification error and averaged over all the task instances, providing the human ($L_{H}$), the AI ($L_{AI}$), and the team performance ($L_{I}$). We also calculated the complementarity potential and effect, including their components (see Section 3). Finally, for each measure, we calculated the average over all the participants.
@@ 4.2.4 Results
Figure 11: Performance results as classification error across conditions, including 95% confidence intervals. The red horizontal line denotes the AI performance.
@@
We observe a significant increase in the inherent complementarity potential (${CP}^{inh}$) in the condition with the complementary AI (Baseline AI: 0.0640, Complementary AI: 0.2480; $d=$ 3.81, $p<$ 0.001, two-tailed Mann-Whitney U test). This reflects a higher level of capability asymmetry between humans and the AI. Compared to the baseline AI condition, the AI makes more erroneous decisions for instances that humans can process correctly, whereas humans err for instances that AI can classify correctly. Conversely, we observe a significant decrease in the collaborative complementarity potential (${CP}^{coll}$) in the condition with the complementary AI (Baseline AI: 0.2026, Complementary AI: 0.0186, $d=$ 3.81, $p<$ 0.001, two-tailed Mann-Whitney U test) as the $CP$ remains constant due to the AI being individually more accurate than the humans. Whereas the inherent complementarity potential constitutes 24% of the overall complementarity potential in the baseline condition, this share rises to 93% in the complementary AI condition. This means that, for the majority of the task instances, one team member is theoretically capable of making a correct decision.
@@
In the baseline condition, 58% of the inherent complementarity potential ($\frac{{CE}^{inh}}{{CP}^{inh}}$) could be realized by the humans integrating their own decision and that of the AI into a final team decision, resulting in a ${CE}^{inh}$ of 0.0368. In the condition with the complementary AI, it is possible to realize 89% of the inherent complementarity potential, resulting in a ${CE}^{inh}$ of 0.2196. This shows a significant performance improvement ($d=$ 3.43, $p<$ 0.001, two-tailed Mann-Whitney U test), which is attributable to a significantly larger fraction of $\frac{{CE}^{inh}}{{CP}^{inh}}$ that could be realized ($d=$ 1.02, $p<$ 0.001, two-tailed Mann-Whitney U test). It indicates that humans tended to rely on the AI decisions when they were correct, but on their decision when it was incorrect. Moreover, in both conditions the collaborative complementarity effect (${CE}^{coll}$) is negative. Whereas the value is only slightly negative in the baseline condition (Baseline AI: -0.0175), it decreases to -0.0990 in the condition with the complementary AI. The difference between the two conditions is significant ($d=$ 1.32, $p<$ 0.001, two-tailed Mann-Whitney U test).
Figure 12: Result summary of the image classification experiment.
The observed changes in both components reveal interesting insights. The increase in the inherent complementarity effect reveals that humans continue to regard the AI as a team member they can rely on, especially with regard to the instances that they find difficult, even though they have observed the AI making several incorrect decisions in respect of the instances that they find easier. Conversely, the decrease in the collaborative complementarity effect reveals that there are also instances where humans, incorrectly, do not rely on the AI. This is an interesting finding, since it could have been conceivable for humans to avoid AI suggestions to a greater extent after witnessing incorrect AI decisions (Dietvorst et al., 2015). Consequently, in the team setting in which the human integrates his/her own decision and that of the AI into a final team decision, there might be a trade-off between turning a higher level of capability asymmetry into performance synergies and not relying on the AI’s suggestions due to witnessing erroneous decisions in “easier” instances. We refer to Appendix C.4 for additional analyses, including an analysis of each image. In this context, similar to the first behavioral experiment, we find a positive collaborative complementarity effect which occurs for individual participants for three task instances (see Table C.4 in Appendix C.4).
In summary, ${CE}^{inh}$ and ${CE}^{coll}$ result in the total complementarity effect ($CE$)—equivalent to the performance difference between the best individual team member and the team performance (Baseline AI: 0.0193, Complementary AI: 0.1206). Figure 12 summarizes the analysis.
@@ 5 Discussion
Our research emphasizes the importance of complementarity and attaining CTP in human-AI decision-making as an important area of human-AI collaboration (Lai et al., 2023). In the past, when early AI-based systems still had limited capabilities, CTP was rarely a focus, as these systems could only provide humans with (partial) decision support within a small range of low-stakes decision-making tasks, e.g., calculating decision-making inputs (Power, 2008). Today, however, AI can perform a growing number of tasks independently with human-level performance and is increasingly utilized in high-stakes decision-making domains such as medicine (McKinney et al., 2020), law (Hillman, 2019), and finance (Day et al., 2018). It can also provide general-purpose support, driven by recent advances in large language models that enable applications like ChatGPT (Bubeck et al., 2023). While this opens up the potential for automation, it also enables a new form of collaboration in which AI and humans are equitable team members.
@@ 5.1 Contributions
Current research lacks a concise concept of human-AI complementarity, as the majority of studies focusing on human-AI collaboration in decision-making do not achieve CTP (Bansal et al., 2021; Hemmer et al., 2021). Moreover, merely observing whether CTP has been achieved or not does not allow in-depth conclusions to be drawn about the synergetic potential of humans and AI through collaboration. In this work, we contribute to a comprehensive understanding of human-AI teams’ inner workings regarding decision-making and provide support on how to achieve CTP more consistently.
@@
More specifically, we develop a conceptualization of human-AI complementarity that introduces and formalizes the notion of complementarity potential and complementarity effect and outlines relevant sources of complementarity. This conceptualization allows for analyzing human-AI teams’ potential for decision-making synergies by making this potential measurable. By differentiating between the inherent and collaborative components of complementarity potential and its realized effect, it is also possible to gain important insights into the collaboration’s functioning in terms of the optimal joint team decision. These insights, along with the sources of complementarity, can inform human-AI teams’ future designs and collaboration mechanisms, depending on the use case and application domain. Furthermore, we not only demonstrate the conceptualization’s utility in two behavioral experiments with humans in the role of integrating the final team decision, but we also empirically show that both sources of complementarity—information and capability asymmetry—can actually lead to CTP. In this context, our conceptualization allows us to reveal interesting empirical insights. The first experiment highlights that providing humans with unique contextual information not only affects the inherent complementarity potential, but can also disproportionally increase the realized amount, i.e., the inherent complementarity effect. This means that the effectiveness of the integration that humans conducted improved, which is an interesting finding. Intuitively, the perception of having more information than the other team member could alternatively result in reduced utilization of the AI suggestions in the team prediction—resulting in a sub-optimal team performance (Jussupow et al., 2020; Mahmud et al., 2022; Sieck and Arkes, 2005). The second experiment reveals that if the human-AI team’s capability asymmetry is greater, this can contribute to the realization of a larger share of inherent complementarity potential. This is also an insightful observation, since humans observing erroneous AI decisions for instances that they could solve relatively easily themselves might lead them to refute AI support even in situations when it is actually helpful (Dietvorst et al., 2015). This finding also contributes to our understanding of team diversity’s compositional impact on performance (Horwitz, 2005). Whereas expertise diversity in human teams could be a performance driver, e.g., because it fosters a broader range of cognitive skills (Cohen and Bailey, 1997), it could also turn into a performance inhibitor, e.g., due to the potential difficulties of achieving a mutual understanding (Dougherty, 1992). As in human teams, there might also be a trade-off, since humans could start avoiding AI suggestions if they observe them to be incorrect too often (Dietvorst et al., 2015). Our conceptualization enables an analysis of this trade-off—finding that, in the specific experiment, the capability asymmetry between humans and AI affects team performance positively. Finally, the results of both experiments provide empirical evidence for the validity of our conceptualization. They empirically support the existence of both the inherent and the collaborative components. In particular, we find that not only are humans able to realize the inherent complementarity potential in the team, but also that the interaction between humans and AI can lead to more accurate team decisions for several task instances compared to their individual decisions, thus proving the existence of the collaborative component.
@@ 5.2 Theoretical Implications
From a theoretical perspective, the proposed conceptualization provides a foundation for future human-AI collaboration research. We offer the research community concrete measures that allow the investigation of human-AI teams’ inner workings at a deeper level than aggregated performance comparisons in behavioral experiments allow. In this context, these measures could also help researchers formulate and test hypotheses about the effect of different socio-technical factors on the realization of decision-making synergies in human-AI teams (Jain et al., 2021).
Our research’s most important implication is the need to design for CTP, which is influenced by the source of complementarity and the collaboration mechanism to derive a team decision. Both should be taken into account purposefully. Increasing unique knowledge in the team positively affects the inherent complementarity potential. From an AI perspective, this could be achieved by deliberately creating “complementary” AI models designed for teamwork that perform well in areas of the feature space where humans experience difficulties (Hemmer et al., 2022; Mozannar and Sontag, 2020; Wilder et al., 2020). From a human perspective, humans could be trained to focus on their unique capabilities and to build awareness to use unique contextual information. Over time, the capabilities of humans and AI may also change. While humans become more experienced, AI models may get updated. For human-AI teams to also function in the long term, it should be ensured that the capabilities of both team members remain complementary to each other over time (Spitzer et al., 2023), which could be monitored by employing the measures of our conceptualization. Lastly, attaining CTP also implies that the collaboration mechanism should be consciously designed to improve the realization of inherent and collaborative complementarity potential.
@@ 5.3 Managerial Implications
Our work also has important implications for managerial decision-makers. In application areas with suitable decision-making tasks, responsible managers should focus on deploying AI systems that enable the realization of CTP through collaboration. If they don’t, their competitors could gain advantages. They could start by collecting data to train AI models that are compatible with teamwork and invest in training their employees in order to enhance their human-AI teams’ inherent complementarity potential.
The conceptualization supports the identification of decision-making use cases suitable for human-AI collaboration by providing the means to measure the complementarity potential and its realized effect. By fostering a deep understanding of the effectiveness of the human-AI team, it enables the targeted improvement of the team’s collaboration. Rather than fearing automation, decision-makers should explore the benefits of working with AI. This is an important perspective, also in the broader societal debate on the role of AI in automation and in the future of work. Our research provides managers with valuable guidance by helping them determine when and how to collaborate with AI to achieve CTP more consistently.
@@ 5.4 Limitations
Our current research has several limitations that future work needs to address. First, the prevalent use of laboratory-based experiments and vignette methodologies in extant literature on human-AI collaboration in decision-making, including our own, could limit our work’s generalizability and its practical implications. This underscores the need for future studies to use a nuanced approach to prevent contributing to the fragmentation of research in this domain. Furthermore, we focus on developing measurement tools, delineating sources of complementarity, and validating the proposed concepts in behavioral experiments. While our experiments’ controlled settings allow us to derive the insights presented in this work, they do not yet address the wide range of complex human factors, such as motivation (Schunk, 1995), engagement (Benz et al., 2024; Chandra et al., 2022), self-efficacy (Westphal et al., 2024), or behavior patterns within teams (Schecter et al., 2022), which also contribute to the effectiveness of human-AI teams (Chandra et al., 2022). In this context, many factors in socio-technical systems are highly interconnected and influence the underlying systems’ acceptance and use (Jain et al., 2021). For example, humans could experience emotional reactions towards the AI system or even the task (Han et al., 2022; Pfeuffer et al., 2019). In addition, situational factors, such as time pressure, are also likely to affect the human-AI team’s ability to realize complementarity potential (Cao et al., 2023; Swaroop et al., 2024). While a broader exploration of these factors is beyond the scope of our work, advancing the field of human-AI collaboration in decision-making further also requires their incorporation as well as finding approaches to manage these factors. We hope that our research will serve as a foundation for studies aiming at expanding knowledge of these mechanisms.
Another limitation is the way in which we measure the counterfactual human decision if the user has not received advice on using AI. In this work, we used a sequential decision-making setup to first measure the human decision and, thereafter, the team decision. However, the sequential nature of the decision-making process could also influence human behavior. Consequently, the timing when the AI’s recommendation is revealed is another critical aspect (Jussupow et al., 2021). Giving the participants the AI’s recommendation upfront could lead to cognitive capacity being less invested in the task because the AI has already provided a possible answer (Green and Chen, 2019).